# Language-Guided Self-Supervised Video Summarization Using Text Semantic Matching Considering the Diversity of the Video

### v1

@@ 3.2. Proposed PDL
Figure 3. Examples of the visualization of diversity of the video. The cosine similarity between adjacent scenes are shown. A lower value indicates a reduced similarity to the adjacent scene.
When a video is linguistically represented through an image captioning model, scenes with little variation tend to generate semantically similar captions continuously. In such instances, uniformly similarity scores are generated, making it challenging to perform effective shot selection. Therefore, we analyze the diversity of the video in accordance with its characteristics to enable appropriate adjustments and mathematically define the diversity of a video represented linguistically. Firstly, based on $E_{C_{i}}$, we employed Kernel Temporal Segmentation (KTS) (Potapov et al., 2014), an algorithm used in conventional video summarization (Zhou et al., 2018; Zhao et al., 2021; Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023) for segmenting frames into shots, to segment the linguistically represented video into scenes of similar content. The number of scenes is denoted as $q$ and the averaged feature value as $E_{C_{i}}$ within the same $j$-th scene as $E_{\rm{scene}_{j}}$. When the $j$-th scene consists of $p$ captions, $E_{\rm{scene}_{j}}$ is calculated as follows:
(8) $E_{\rm{scene}_{j}}=\frac{1}{p}\sum_{i=1}^{p}E_{C_{i}}.$
When the cosine similarity between $j$-th scene and ($j+1$)-th scenes is $k$-th adjacent scene, we define that as $s_{\rm{change}_{k}}$ and calculated as follows:
Figure 3 is the examples of the transition of $s_{\rm{change}_{k}}$. It shows that in adjacent scenes, $s_{\rm{change}_{k}}$ is higher for visually similar videos. This is attributed to the generation of semantically similar captions for scenes that resemble each other. As a result, when the cosine similarity between these captions is calculated, the score approaches 1. Then, the average value of $s_{\rm{change}_{k}}$ is defined as $D$ and is calculated as follows :
(10) $D=\frac{1}{q-1}\sum_{j=1}^{q-1}s_{\rm{change}_{k}}.$
Figure 4 shows the distribution of $D$ in videos within video summarization dataset: SumMe (Gygli et al., 2014) and TVSum (Song et al., 2015). It shows that videos in the SumMe dataset have high $D$ scores, indicating that there are many semantically similar scenes. On the other hand, videos in the TVSum dataset have relatively many low $D$ scores, meaning there is significant variation between adjacent scenes.
@@
Figure 4. The histogram of $D$ for individual videos within SumMe (Gygli et al., 2014) and TVSum (Song et al., 2015) datasets. A lower value indicates that the video has greater diversity.
(11) $\lambda=\begin{cases}0&\text{if }D<\delta\\ De^{D}&\text{if }D\geq\delta,\end{cases}$
where $\delta$ is a threshold to measure the diversity of the video. The values of 0 and $De^{D}$ are decided empirically.
@@ 4.1. Datasets
We evaluate our method on two standard video summarization datasets, SumMe (Gygli et al., 2014) and TVSum (Song et al., 2015), to compare with previous works. SumMe comprises 25 unedited personal YouTube videos that capture various events, such as cooking and sports, with each video ranging in length from 30 seconds to 6 minutes. The title of the video is available as metadata. Annotations were created over a total of more than 40 hours by 15 to 18 annotators.
TVSum contains 50 edited YouTube videos, spanning 10 categories including dog shows and parades, with 5 videos from each category. Each video lasts between 1 to 10 minutes. Its metadata encompasses titles, genres, and query categories. The annotations are provided by 20 annotators.
@@ 4.2. Implementation Details
Following existing works (Zhou et al., 2018; Mahasseni et al., 2017; Jung et al., 2020; Zhang et al., 2016; Zhao et al., 2021; Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023), we first downsample the input video to two frames per second (fps) and process the individual frames with pre-trained GIT (Wang et al., 2022). When generating individual captions, we use “a photo of” as a prompt and the prompt is excluded when generating text summary. One main reason for using an image captioning model instead of a video captioning model for caption generation is that although video captioning models can input sequences of frames, using an image captioning model allows for more accurately assessment of the content of each frame and calculation of frame-level scores. Regarding the diversity of captions for each frame, 69% of captions in the SumMe dataset (Gygli et al., 2014) and 76% of captions in the TVSum dataset (Song et al., 2015) describe differently from other captions, indicating that there was little duplication in the captions for each frame.
Then, we obtain 1024-dimensional output from the pool5 layer of GoogLeNet (Szegedy et al., 2015) pre-trained on ImageNet (Russakovsky et al., 2009) for the sampled frames. Using those frame-level features, each video frame is segmented into several shots by KTS algorithm (Potapov et al., 2014) as in previous video summarization task.
When generating text summaries, we use the Chain of Density prompt that incorporates each video’s metadata and a description of the task, based on the prompt proposed by Adams et al. (Adams et al., 2023). We train our model with a learning rate of $5\times 10^{-5}$ and $1\times 10^{-5}$, the margin $m$ in Eq. (4) of 0.11 and 0.06 for the SumMe and TVSum datasets, respectively. On both datasets, our model is optimized by Adam and maximum training epoch is 100. The threshold $\delta$ defined in Eq. (10) is used to quantify the diversity of the input video and is empirically set 0.65. The number of videos where $D$ exceeds 0.65 is 14 for SumMe and 3 for TVSum. After calculating frame-level scores, the conversion of frame-level scores to shot-level scores is the same as previous works (Zhou et al., 2018; Zhao et al., 2021; Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023). These frame-level scores are aggregated into shot-level importance scores by averaging the scores within each shot. Then, important shots are selected by solving the Knapsack problem.
@@ 4.3. Evaluation Metrics
F-score. In order to compare with other state-of-the-art (SOTA) unsupervised video summarization methods, we use F-score as the evaluation metric. It measures the overlap between the predicted video summary and the reference human summary. Specifically, given the automated video summary $A$ and the human summary $H$, the precision $p$ and recall $r$ are defined and F-score is computed as follows:
(12) $p=\frac{A\cap H}{\text{Len}(A)},\quad r=\frac{A\cap H}{\text{Len}(H)},\quad F=\frac{2pr}{p+r}$
Rank-based Evaluation. The F-score are found to be significantly influenced by the pre-processing segmentation process and even a random method is able to reach similar scores (Otani et al., 2019). To address this issue, two rank-based evaluation, Kendall’s $\tau$ and Spearman’s $\rho$, are used to evaluate the quality of the model. In these metrics, the predicted frame-level scores are compared with the scores annotated by humans.
@@ 4.4. Results
Table 1. Comparison with SOTA unsupervised learning methods on the SumMe (Gygli et al., 2014) and TVSum (Song et al., 2015) datasets using F-score (F), Kendall’s $\tau$ and Spearman’s $\rho$ metrics. The bolded and underlined items represent the optimal and sub-optimal results.
SumMe TVSum
Methods F $\uparrow$ $\tau\uparrow$ $\rho\uparrow$ F $\uparrow$ $\tau\uparrow$ $\rho\uparrow$
Random (Otani et al., 2019) 0.410 0.000 0.000 0.570 0.000 0.000
Human (Otani et al., 2019) 0.540 0.205 0.213 0.540 0.177 0.204
DRDSN (Zhou et al., 2018) 0.414 0.047 0.048 0.576 0.020 0.026
CSNet (Jung et al., 2019) 0.513 - - 0.588 - -
$\text{RSGN}_{u}$ (Zhao et al., 2021) 0.423 0.071 0.073 0.580 0.048 0.052
DSAVS (Zhong et al., 2022) 0.466 - - 0.591 0.080 0.087
CASUM (Apostolidis et al., 2022) 0.511 - - 0.614 0.160 0.210
RSSUM (Abbasi and Saeedi, 2023) 0.525 - - 0.614 - -
Ours 0.404 0.102 0.138 0.590 0.136 0.178
We compared our proposed method with the SOTA unsupervised learning video summarization methods on the SumMe and TVSum datasets. The results are shown in Table 1. As indicated in Table 1, our method surpasses most of the compared SOTA methods, except for CASUM (Apostolidis et al., 2022) and RSSUM (Abbasi and Saeedi, 2023). However, it should be noted that CSNet, DSAVS, CASUM, and RSSUM use self-attention mechanisms (Vaswani et al., 2017; Vyas et al., 2020), which suffer from high computational complexity that escalates with increased input video length, presenting scalability challenges for processing longer videos. In contrast, our method calculates semantic textual similarity for individual captions and the text summary separately, enabling calculations independent of video length, with the exception of score averaging. It ensures consistent performance across videos of any length, making it well-suited for diverse real-world applications.
@@ 4.5. Model Analysis
SumMe TVSum
Method Loss $\tau\uparrow$ $\rho\uparrow$ $\tau\uparrow$ $\rho\uparrow$
Margin ranking loss only $\mathcal{L}_{m}$0.015 -0.019 0.136 0.178
@@
Unweighted sum of losses $\mathcal{L}_{m}+\mathcal{L}_{s}$ 0.037 0.051 0.024 0.031
2 task uncertainty weighting (Kendall et al., 2018) $\frac{1}{2\sigma_{1}^{2}}\mathcal{L}_{m}+\frac{1}{2\sigma_{2}^{2}}\mathcal{L}_{s}+\text{log}\sigma_{1}+\text{log}\sigma_{2}$ 0.037 0.051 0.024 0.031
PDL (Ours) $\mathcal{L}_{m}+\lambda\mathcal{L}_{s}$ 0.102 0.138 0.136 0.178
Our proposed PDL is comprised of two components: the margin ranking loss (${L}_{m}$) and regularization loss (${L}_{s}$(Mahasseni et al., 2017). The effectiveness of our PDL is demonstrated in Table 2, where it consistently achieves the highest scores across all evaluation metrics on both datasets. For a fair comparison, we use the same experimental settings such as $m$ and learning rate. In our analysis, we assess the impact of each loss individually, as well as their combined effect through simple addition with equal weights, and through the balanced parameter approach we propose. Additionally, we explore a technique proposed by Kendall et al (Kendall et al., 2018) for combining loss functions in multi-task learning frameworks which can learn a relative weighting automatically from the data. In our settings, margin ranking loss ${L}_{m}$ addresses the main task, while regularization loss ${L}_{s}$ is tailored for the secondary task. This method leverages the concept of homoscedastic uncertainty, which provides a measure of the inherent noise associated with each task. By using this uncertainty as a weighting factor, the loss function is able to adaptively prioritize tasks during training, reflecting their relative difficulty and importance. In this case $\sigma_{1}$ and $\sigma_{2}$ written in Table 2 act as observation noise scalers. This approach enables the network to learn the optimal balance between different tasks, accounting for the fact that not all tasks contribute equally to the performance of the model.
The results suggest that employing margin ranking loss exclusively to optimize our model proves effective for the TVSum dataset, which features videos with numerous scene changes, but not to the same extent for the SumMe dataset. Conversely, only using regularization loss is more effective than optimizing with only margin ranking loss in SumMe dataset but not in the TVSum dataset. Both the unweighted sum of losses, ${L}_{m}+{L}_{s}$, which treats both tasks with equal importance and the automatic weighted loss proposed by Kendall et al. fail to account for the relative contributions and importance of each task on both datasets. On the other hand, our proposed PDL, ${L}_{m}+\lambda{L}_{s}$, which uses balanced parameters that take into account the diversity of the video, allows the model to dynamically prioritize between tasks during training based on the data, resulting in the most robust and effective model. Therefore, we not only demonstrate the superior capabilities of our PDL but also elucidate the significance of a tailored, data-driven approach to loss function formulation.
@@ 4.6. Visualization
Figure 5. Visualization results of the summarized video 6 in SumMe (“Car railcrossing”) generated by our models using different self-supervised loss function. $D$ score of the video is 0.617. The light-gray bars in the figure represent the ground truth importance scores, while the orange areas indicate the parts selected by the model. The x-axis represents the frame index. The black vertical lines in the figure represent significant content changes within this video, with details documented in this section. The five images below are the representative frames selected as the video summary.
@@
Figure 6. Visualization results of the summarized video 11 in TVSum (“Pet Joy Spa Grooming Services - Brentwood, CA”) generated by our models using different self-supervised loss function. $D$ score of the video is 0.713. The light-gray bars in the figure represent the ground truth importance scores, while the orange areas indicate the parts selected by the model. The x-axis represents the frame index. The five images below are the representative frames selected as the video summary.
@@
The original video 6 in SumMe depicts a vehicle encountering an unexpected obstacle while crossing a railroad track, resulting in a collision. Following the incident, the video showcases a collaborative effort involving a backhoe loader and individuals working together to assist and recover the vehicle. In the Figure  5, the black vertical line delineates scenes before and after the vehicle collides with the obstacle. This line signifies a major branching point that significantly divides the video content into two parts, and extracting summary scenes before and after this point holds qualitatively significant meaning. The analysis of the result shown in Figure 5 (a) reveals that the comparative method only includes scenes after the collision in the summarized video. In contrast, our proposed method shown in Figure 5 (b) successfully captures scenes both before and after the collision, demonstrating its ability to encapsulate the entire storyline. This distinction underscores the effectiveness of our approach in providing a more comprehensive and contextually rich video summary.
The original video 11 in TVSum revolves around the facilities of a pet spa shop and Figure 6 shows the visualization of the results. Despite the minimal scene changes within the original video, the video summary generated by our proposed method is more diverse, successfully picking up qualitatively important parts from a wide range of content. Moreover, as almost all of the peak areas of the ground truth are selected, our proposed method demonstrates a significantly higher F-score, showcasing its utility and effectiveness in creating meaningful summaries even from footage with limited variation in scenes. This indicates the proposed method’s capability to identify and include key moments, ensuring a comprehensive and engaging summary.

### v2

@@ 3.3. Impact of the loss functions
Figure 4. The histogram of $D$ for individual videos within the SumMe (Gygli et al., 2014) and TVSum (Song et al., 2015) datasets. A lower value indicates that the linguistically represented video has lower diversity. TVSum has more videos with higher $D$ scores compared to SumMe.
The improved margin ranking loss focuses on enlarging score difference between frame captions and the text summary, and the regularization loss aims at control the sparsity of the selection vector (Mahasseni et al., 2017). We investigated the contributions of the regularization loss. We conducted experiments, assigning a fixed value ($\alpha$) to the regularization loss. The overall loss was constructed as follows:
(6) $\mathcal{L}=\mathcal{L}_{m}+\alpha\mathcal{L}_{s}.$
Figure 4 shows the quality of the generated video summaries when varying $\alpha$ using two video summarization datasets: SumMe (Gygli et al., 2014) and TVSum (Song et al., 2015). It shows that a larger $\alpha$ value in SumMe results in higher quality video summaries, while in TVSum, a smaller $\alpha$ produces better quality.
3.4. Proposed PDL
@@
Figure 5. Examples of the transition of $s_{\rm{change}_{k}}$. The cosine similarity between adjacent scenes are shown. A lower value indicates a reduced similarity to the linguistically represented adjacent scene.
Based on the aforementioned observations, the $\alpha$ value should be defined differently to account for the varying characteristics of different datasets. An adaptive loss function is required to effectively handle general video summarization tasks. We also find that these differences are fundamentally linked to the diversity of the videos within each dataset. Therefore, we investigate the diversity of the datasets. Firstly, based on encoded captions generated from all the video frames, we employed Kernel Temporal Segmentation (KTS) (Potapov et al., 2014), an algorithm used in conventional video summarization (Zhou et al., 2018; Zhao et al., 2021; Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023) for segmenting frames into scenes, to segment the linguistically represented video into scenes of similar content. The number of scenes is denoted as $q$, and the averaged feature value of the encoded captions generated from all the video frames within the same $j$-th scene is represented as $E_{\rm{scene}_{j}}$. When the $j$-th scene consists of $p$ video frames, $E_{\rm{scene}_{j}}$ is calculated as follows:
(7) $E_{\rm{scene}_{j}}=\frac{1}{p}\sum_{i=1}^{p}E_{C_{i}}.$
When the cosine similarity between $j$-th and ($j+1$)-th scenes is $k$-th adjacent scene, we define it as $s_{\rm{change}_{k}}$ and calculated as follows:
Then, the similarity between the linguistically represented video frames is defined as ${\rm{sim}}_{\rm{scene}}$, and $D$ is defined as the diversity of the video, calculated as follows:
(9) ${\rm sim}_{{\rm scene}}=\frac{1}{q-1}\sum_{j=1}^{q-1}s_{\rm{change}_{k}},\quad D=1-{\rm sim}_{{\rm scene}}.$
Note that a higher ${\rm sim}_{{\rm scene}}$ results in a lower $D$ score, indicating that when the similarity between adjacent scenes is high, the diversity of the video is low. Figure 4 shows the distribution of $D$ in videos within SumMe and TVSum. It shows that videos in SumMe have lower $D$ scores, suggesting the presence of many semantically similar scenes. Conversely, videos in TVSum have relatively higher $D$ scores, indicating the diversity of each video in the dataset is high. Also, Figure 5 shows examples of the transition of $s_{\rm{change}_{k}}$. In adjacent scenes, $s_{\rm{change}_{k}}$ is higher for linguistically similar videos. In contrast, linguistically different videos have lower $s_{\rm{change}_{k}}$
Building upon the aforementioned findings, we introduce a novel PDL, denoted as $\mathcal{L}_{\rm PDL}$. It adapts the contribution of the regularization loss based on video diversity and can effectively function across different video domains. The mathematical definition of $\mathcal{L}_{\rm PDL}$ is
(10) $\mathcal{L}_{\rm PDL}=\mathcal{L}_{m}+\lambda\mathcal{L}_{s},$
where $\lambda$ is an adaptive value to dynamically adjust the contribution of $\mathcal{L}_{s}$, determined by the video diversity and defined as follows:
(11) $\lambda=\begin{cases}0&\text{if }D\geq\delta,\\ (1-D)\exp(1-D)&\text{if }D<\delta,\end{cases}$
where $\delta$ denotes a threshold to measure the diversity of the video. The values of 0 and $(1-D)\exp(1-D)$ are decided empirically. When the video diversity is high, incorporating the regularization loss becomes unnecessary, as inherent diversity ensures a rich and comprehensive representation of the content. Conversely, if similar captions are prevalent, making it challenging to differentiate based on similarity, it is necessary to reduce the contribution of the improved margin ranking loss.
@@ 4.1. Datasets
We evaluate our method on two standard video summarization datasets, SumMe (Gygli et al., 2014) and TVSum (Song et al., 2015), to compare it with previous works (Zhou et al., 2018; Jung et al., 2019; Zhao et al., 2021; Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023).
SumMe comprises 25 unedited personal YouTube videos that capture various events, such as cooking and sports, with each video ranging in length from 30 seconds to 6 minutes. The title of the video is available as metadata. Annotations were created over a total of more than 40 hours by 15 to 18 annotators, with the audio track not included.
TVSum contains 50 edited YouTube videos, spanning 10 categories including dog shows and parades, with 5 videos from each category. Each video lasts between 1 to 10 minutes. Its metadata encompasses titles, genres, and query categories. The annotations are provided by 20 annotators, who watched the videos without audio.
@@ 4.2. Implementation Details
We first downsample the input video to two frames per second as existing works (Zhou et al., 2018; Jung et al., 2019; Zhao et al., 2021; Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023). Then, we process the downsampled frames with pre-trained GIT (Wang et al., 2022) using “a photo of” as a prompt and the prompt is excluded when generating the text summary. For the text summaries, we use the Chain of Density prompt that incorporates each video’s metadata, based on the prompt proposed by Adams et al. (Adams et al., 2023). On both datasets, our model is optimized by the Adam optimizer, with a maximum training epoch of 100. The threshold $\delta$ defined in Eq. (9), to quantify the diversity of the input video is established at 0.35. The number of videos with $D$ below 0.35 is 14 for SumMe and 3 for TVSum. To convert frame-level scores into scene-level scores, we segment the video into scenes as previous works (Zhou et al., 2018; Jung et al., 2019; Zhao et al., 2021; Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023). First, we extract 1024-dimensional features from the pool5 layer of GoogLeNet (Szegedy et al., 2015) pre-trained on ImageNet (Russakovsky et al., 2009) for the video frames. Using these frame-level features, the video is segmented into several scenes using the KTS algorithm (Potapov et al., 2014). After calculating frame-level importance scores, these scores are aggregated into scene-level importance scores by averaging the scores within each scene. Finally, important scenes are selected by solving the knapsack problem, ensuring the video summary is 15% of the original video’s length, a common method for video summarization (Zhou et al., 2018; Jung et al., 2019; Mahasseni et al., 2017; Jung et al., 2020; Zhang et al., 2016; Zhao et al., 2021; Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023; Wang et al., 2024; He et al., 2023; Li et al., 2023a; Ghauri et al., 2021; Argaw et al., 2024; Tian et al., 2024; Jiang and Mu, 2022; Narasimhan et al., 2021; Son et al., 2024). More details about the experimental settings are provided in the supplementary material.
@@ 4.3. Evaluation Metrics
When comparing our method with other state-of-the-art (SOTA) unsupervised video summarization methods, we omit the F-score, which measures the overlap between the predicted video summary and the reference summaries. This choice is based on research papers that suggest the F-score is unreliable in the video summarization task (Otani et al., 2019; Terbouche et al., 2023). The F-score is influenced by the common segmentation and segment selection process. Otani et al. (Otani et al., 2019) demonstrated that even randomly selecting these pre-processed segments can achieve high F-scores. Also, when solving the knapsack problem, the model selects as many short scenes as possible instead of choosing a longer scene with higher scene-level score. Therefore, we use two rank-based evaluations, Kendall’s $\tau$ (Kendall, 1945) and Spearman’s $\rho$ (Zwillinger and Kokoska, 1999), proposed in (Otani et al., 2019), as evaluation metrics. In these metrics, the predicted frame-level scores are compared with the scores annotated by humans, which are independent of the segmentation and segment selection process.
@@ 4.4. Results
Table 1. Comparison with SOTA unsupervised learning methods on the SumMe and TVSum datasets using Kendall’s $\tau$ and Spearman’s $\rho$ metrics. The bolded and underlined items represent the best and second-best results.
Methods SumMe TVSum
$\tau\uparrow$ $\rho\uparrow$ $\tau\uparrow$ $\rho\uparrow$
Random (Otani et al., 2019) 0.000 0.000 0.000 0.000
Human (Otani et al., 2019) 0.205 0.213 0.177 0.204
DRDSN (Zhou et al., 2018) 0.047 0.048 0.020 0.026
CSNet (Jung et al., 2019) - - 0.025 0.034
$\text{RSGN}_{u}$ (Zhao et al., 2021) 0.071 0.073 0.048 0.052
DSAVS (Zhong et al., 2022) - - 0.080 0.087
CASUM (Apostolidis et al., 2022) 0.063 0.084 0.160 0.210
RSSUM (Abbasi and Saeedi, 2023) 0.007 0.015 0.080 0.106
Ours 0.102 0.138 0.133 0.174
We compared our proposed method with the SOTA unsupervised learning video summarization methods on the SumMe and TVSum datasets. The results are shown in Table 1, and our method achieves the best performances on SumMe and the second-best results on TVSum. CASUM shows the best performance on TVSum but it uses different $\epsilon$ values in Eq. (5) for videos within the same dataset, which is significantly different from others that use the same hyperparameters for all videos within each dataset. Moreover, our proposed method has another advantage in terms of computational cost. DSAVS and CASUM utilize Query-Key-Value attention mechanisms, and RSSUM utilizes multi-head attention mechanisms (Vaswani et al., 2017; Vyas et al., 2020), which have a $\mathcal{O}(d\cdot n^{2})$ computational complexity for $n$ frames and hidden dimension $d$, growing quadratically with the number of frames $n$. In contrast, ours has a $\mathcal{O}(d\cdot n)$ computational complexity. This scales linearly with the video length, making it more efficient for processing longer videos.
@@ 4.5. Model Analysis
Method Loss SumMe TVSum
$\tau\uparrow$ $\rho\uparrow$ $\tau\uparrow$ $\rho\uparrow$
Improved margin ranking loss only $\mathcal{L}_{m}$0.015 -0.019 0.121 0.159
@@
AWL (Kendall et al., 2018) $\frac{1}{2\sigma_{1}^{2}}\mathcal{L}_{m}+\frac{1}{2\sigma_{2}^{2}}\mathcal{L}_{s}+\text{log}\sigma_{1}+\text{log}\sigma_{2}$ 0.037 0.051 0.024 0.031
PDL (Ours) $\mathcal{L}_{m}+\lambda\mathcal{L}_{s}$ 0.102 0.138 0.133 0.174
Our proposed PDL is comprised of two components: the improved margin ranking loss (${L}_{m}$) and regularization loss (${L}_{s}$(Mahasseni et al., 2017). The effectiveness of our PDL is demonstrated in Table 2, where it consistently achieves the highest scores on both datasets. In our analysis, we assess the impact of each loss individually and explore the combination of loss functions in multi-task learning frameworks, which can learn relative weighting automatically from the data (Kendall et al., 2018). In Table 2, $\sigma_{1}$ and $\sigma_{2}$ act as observation noise scalers. For a fair comparison, we use the same experimental settings.
The results suggest that using only the improved margin ranking loss effectively optimizes the model for the TVSum dataset, which features numerous scene changes. Conversely, using only regularization loss is more effective for the SumMe dataset than the improved margin ranking loss alone, as it helps maximize the diversity in the generated summaries. The Automatic Weighted Loss (AWL) (Kendall et al., 2018) fails to account for the relative contributions and importance of each task in both datasets. On the other hand, our proposed PDL, which uses balanced parameters that take into account the diversity of the video, allows the model to dynamically prioritize between tasks, resulting in the most robust and effective model. Therefore, we not only demonstrate the superior capabilities of our PDL but also highlight the significance of a tailored approach to loss function formulation that specifically considers the diversity of the video in the video summarization tasks.
@@ 4.6. Visualization
Figure 6. Visualization results of the summarized video 6 in SumMe (“Car railcrossing”) generated by our models using different self-supervised loss functions. $D$ score of the video is 0.383. The light-gray bars in the figure represent the ground truth importance scores, while the orange areas indicate the parts selected by the model. The x-axis represents the frame index. The black vertical lines in the figure represent significant content changes within this video, with details documented in this section. The five images below are the representative frames selected as the video summary.
@@
Figure 7. Visualization results of the summarized video 11 in TVSum (“Pet Joy Spa Grooming Services - Brentwood, CA”) generated by our models using different self-supervised loss function. $D$ score of the video is 0.287.
@@
The original video 6 in SumMe depicts a vehicle encountering an unexpected obstacle while crossing a railroad track, resulting in a collision. Following the incident, the video showcases a collaborative effort involving a backhoe loader and individuals working together to assist and recover the vehicle. In Figure  6, the black vertical line delineates scenes before and after the collision, which divides the video content into two parts. The analysis of the result shown in Figure 6(a) reveals that the comparative method only includes scenes after the collision in the summarized video. In contrast, our proposed method shown in Figure 6(b) successfully captures scenes both before and after the collision, demonstrating its ability to encapsulate the entire storyline. This distinction underscores the effectiveness of our approach in providing a more comprehensive and contextually rich video summary.
The original video 11 in TVSum revolves around the facilities of a pet spa shop. Figure 7 shows that despite the minimal scene changes within the original video, the video summary generated by our proposed method is more diverse, successfully picking up qualitatively important parts from a wide range of content.
Moreover, our proposed method demonstrates significantly higher $\tau$ and $\rho$, showcasing its utility and effectiveness in creating meaningful summaries. This indicates the capability to identify and include key moments, ensuring a comprehensive and engaging summary.